# Reasoning Beyond Bias: A Study on Counterfactual Prompting and Chain of Thought Reasoning

### v2

@@ Introduction
Finally, we evaluated the effect of APriCoT on the accuracy of the model. We expected that the accuracy of the model should improve due to a combination of noise reduction from BRP mitigation and a more sophisticated deliberative mechanism. We measure accuracy only across the entire dataset by taking the number of questions answered correctly divided by the total number of questions. The result is an accuracy of $37.96\%$ for CF. Likely due to its susceptibility to BRP effects, CF+CoT saw a slight reduction to $35.56\%$, while APriCoT improved to $42.79\%$.
@@
This alone suggests that APriCoT improves model performance in addition to mitigating surface level biases, but the results improve further with slight decomposition, as shown in Figure 4. When tracked per-question, we see that for every question that was answered correctly with CF+CoT, APriCoT is more likely to also be correct and less likely to be incorrect compared to CF+CoT. To an even greater degree, questions answered incorrectly without CF+CoT are more likely to be answered correct and less likely to be answered incorrectly using APriCoT than CF+CoT. In short, APriCoT improves more than CF+CoT on performance in terms of overall accuracy in every way.
APriCoT shows evidence of being both a less biased and more effective method of reasoning in LLMs compared to the standard practice of CoT.
Prior Work
In this section, we briefly survey the existing literature surround mitigation of base rate effects and the use of CoT in connection with other heuristics and biases. BRP effects were evaluated in some depth by Wei et al. (2024); Moore et al. (2024). Both works found that LLMs are prone to BRP effects in the context of MCQA tasks. Moore et al. (2024) proposed a method for reformulating MCQA tasks to be less sensitive in to BRP effects, both helpful and detrimental, but they do not work to mitigate the effects in model behavior. Wei et al. (2024) attempted to mitigate the BRP effect in behavior by calibrating the output probabilities using the measured BRPs. Chua et al. (2024) attempted to mitigate LLM initial behavior biases in a CoT context using fine-tuning, but the biases being mitigated were user-induced rather than inherent to the model, a distinguishing feature of BRPs.
Prior research has also identified similar undesirable behaviors with CoT. Shaikh et al. (2023) found that CoT is prone to reveal and exacerbate the social biases that the model learned during pre-training. This hints at a consistent bias reinforcement behavior across bias dimensions. Saparov and He (2023) investigated CoT’s utility for complex reasoning and found that models using CoT are capable of strong instantaneous deductive steps but rarely explore multiple semantic paths for more deliberative reasoning, which we attempt to induce explicitly with APriCoT.
@@ Conclusion
In this work, we have attempted to mitigate a simple, but pernicious, bias in LLM behavior that complicates research into other behaviors and abilities. This bias, known as base-rate bias, was shown both here and in previous work to strongly predict model behavior on the MMLU benchmark, which is a common measure of LLM factual knowledge and language understanding. We showed that induction of more complex reasoning via the commonly employed CoT prompting strategy, surprisingly exacerbates the BRP bias, which we argue is the result of confirmation bias-like behavior.
We proposed and evaluated a novel prompting and evaluation method called Agnostically Primed Chain-of-Thought, otherwise known as APriCoT. The method induces more complex and focused deliberation, closer approximating System-2 thought than CoT alone. We showed that APriCoT both significantly reduces, though still does not fully eliminate, BRP effects and noticeably improves model performance on the MMLU task. We hope for this method to be a useful step toward controlling for or eliminating System-1 heuristics in LLM behavior research and that it might inspire useful methods in LLM-based complex reasoning systems.
@@
Our methods were limited to a relatively rigid and strictly defined problem structure. Future work should investigate whether APriCoT is appropriate for use in more complex reasoning tasks like search or planning. In addition, APriCoT, despite being influenced by CF prompting in its creation, makes no assumptions about the evaluation method used. As such, future work should investigate whether APriCoT yields similar improvements for cloze testing and any other alternative instantaneous evaluation techniques.
APriCoT is currently only defined for tasks with a predefined set of available options. Future work may investigate methods that do not require pre-defined options, such as dynamic option generation by the model itself. This could involve complex agentic planning tasks or open ended generation benchmarks.
This work also introduces or highlights multiple theoretical questions regarding mechanistic explanations for LLM behavior. These include the mere existence of anti-confirmation bias, the dependence of CF overt behavior on cloze BRPs, and the lack of correlation between CF overt behavior and CF BRPs. Future theoretical work should try to account for these results.
@@
All experiments described were performed exclusively on the newly released LLaMa 3.1 model from Meta AI (Dubey et al. 2024). Due to resource constraints, only the 8 billion parameter model is evaluated. This model is newly released at the time of writing and has shown promising results on various benchmarks, including state-of-the-art performance on the MMLU benchmark among open source LLMs. We cannot, however, make strong claims of whether the results reported herein will transfer across model families or persist with increased model size.
@@
All experiments were performed with 100 hours on a Google Colab NVIDIA A100 GPU. All code and data relevant to this work are available open source at [GITHUB-LINK-REDACTED] to accommodate replication.

### v3

@@ Introduction
Finally, we evaluated the effect of APriCoT on the accuracy of the model. We expected that the accuracy of the model should improve due to a combination of noise reduction from BRP mitigation and a more sophisticated deliberative mechanism. The result is an accuracy of $37.96\%$ for CF. Likely due to its susceptibility to BRP effects, CF+CoT saw a slight reduction to $35.56\%$, while APriCoT improved to $42.79\%$.
It should be noted that all results reported herein show substantially reduced accuracy on MMLU than was reported at the release of LLaMa 3.1 \@BBOPcitep\@BAP\@BBN(Dubey et al., 2024)\@BBCP. This likely stems from differences in evaluation methodology. We do not have access to the specific methodology used to evaluate LLaMa 3.1. Further, previous work has shown conclusively that subtle changes such as CF prompting in lieu of cloze testing can affect model accuracy - a weakness still present in current LLMs \@BBOPcitep\@BAP\@BBN(Moore et al., 2024)\@BBCP. The seemingly contradictory results between our work and that of \@BBOPcitet\@BAP\@BBNDubey et al. (2024)\@BBCP is evidence that the models and the behaviors they exhibit are complex and multifaceted. The difficulty of isolating behaviors is one of the primary motivations behind this work.
Figure 3: Flow analysis of how CF+CoT and APriCoT each changes CF performance on the MMLU. APriCoT has the highest accuracy at 43% with CF+CoT = 36% and CF = 38%.
@@
While APriCoT improves model performance in addition to mitigating surface level biases, the results improve further with slight decomposition, as shown in Figure 3. When tracked per-question, we see that for every question that was answered correctly with CF, APriCoT is more likely to also be correct and less likely to be incorrect compared to CF+CoT. To an even greater degree, questions answered incorrectly with CF are more likely to be answered correct and less likely to be answered incorrectly using APriCoT than CF+CoT.
@@ Conclusion
In this work, we have attempted to mitigate a simple, but pernicious, bias in LLM behavior that complicates research into other behaviors and abilities. This bias, known as base-rate bias, was shown both here and in previous work to strongly predict model behavior on the MMLU benchmark, which is a common measure of LLM factual knowledge and language understanding. We showed that induction of more complex reasoning via the commonly employed CoT prompting strategy surprisingly exacerbates the BRP bias, which we argue is the result of confirmation bias-like behavior that approximates thinking fast.
We proposed and evaluated a novel prompting method called Agnostically Primed Chain-of-Thought, otherwise known as APriCoT. The method induces more deliberation, more closely approximating slow thinking. Specifically, we showed that APriCoT both significantly reduces, though still does not fully eliminate, BRP effects and noticeably improves model performance on the MMLU task. We hope for this method to be a useful step toward mitigating unintended heuristics in LLM behavior research and that it might inspire useful methods in LLM-based complex reasoning systems.
@@
Future work should investigate whether APriCoT is appropriate for use in more complex reasoning tasks like search or planning. In addition, APriCoT, despite being influenced by CF prompting in its creation, makes no assumptions about the evaluation method used. As such, future work should investigate whether APriCoT yields similar improvements for cloze testing and other evaluation methods.
APriCoT is currently only defined for tasks with a closed set of options. Future work may investigate methods that permit open ended questions, such as dynamic option generation by the model itself.
This work also introduces or highlights multiple theoretical questions regarding mechanistic explanations for LLM behavior that warrant further study. These include confirmation bias, the dependence of CF overt behavior on cloze BRPs, and further work evaluating the nature of LLM reasoning.
Potential Experimental Bias
In this work, we strove to maintain a valid, robust methodology at all stages. We recognize, however, that biases and assumptions we hold may affect the work herein and limit the generalizability of any results reported. In this section, we identify and address those biases of which we are aware to aid in the application and extension of this research.
This work is limited to experiments on a specific model and dataset, each chosen with care. The model, Meta LLaMa 3.1 8B \@BBOPcitep\@BAP\@BBN(Dubey et al., 2024)\@BBCP, was the state of the art in small scale open-weight LLMs at experimentation time. We limited experiments to open-weight models to maximize reproducibility. Resource constraints limited experimentation to the smaller 8B variant. While we recognize that the results reported herein may not generalize to larger variants, we believe that they represent important findings and hope that these results inspire further study on a wider variety of models. Similarly, the MMLU benchmark is a mainstay in the evaluation of new LLMs at release. We believe this makes any results concerning MMLU’s efficacy and methods for ensuring its efficacy to be of importance to the research community.
@@
APriCoT method is notably computationally expensive, increasing inference costs by a factor equal to the number of choices presented (4 in the case of MMLU). We stress that this method is intended only for model evaluation, not end-user applications. We believe that a one-time increase in cost of evaluation is justified in exchange for a more trustworthy and transparent gauge of model capabilities. We also hope that future work may provide similar benefits without the increase in computational expense.
@@
All experiments were performed with $\sim$100 hours on a Google Colab NVIDIA A100 GPU. All code and data are available at www.github.com/KyleAMoore/APriCoT_Experiments.